OpenStudy (nuccioreggie):

Solve the system of equations and choose the correct answer from the list of options. x + y = −3 y = 2x + 2 five over 3 comma 4 over 3 negative 5 over 3 comma negative 4 over 3 negative 3 over 5 comma negative 3 over 4 3 over 4 comma 3 over 5

OpenStudy (anonymous):

Ummm...

OpenStudy (anonymous):

It is 3/4,3/5

OpenStudy (anonymous):

Another medal! XD

OpenStudy (nuccioreggie):

EXPLAIN HOW

OpenStudy (nuccioreggie):

@Sushi121212 help is he / her correct

OpenStudy (anonymous):

Ok so lets move the y over to the right which gives us -3y and _/-3y is equal to .75, which is eqaul to 3/4

OpenStudy (anonymous):

2x + 2 = 4x and _/4x = 0.60 which is equal to 3/5

OpenStudy (anonymous):

medal pweez :3

OpenStudy (sushi121212):

The answer is negative 5/3 and 4/3 because: x + y = −3 y = 2x + 2 -5/3+-4/3=-3 (see below for work) 5 4 9 - ___ + - ___ = -__ or -3 3 3 3 -4/3 = 2(-5/3)+2 -4/3 = -10/3+2 -4/3 = -4/3

OpenStudy (anonymous):

@Sushi121212 could u help me ?

OpenStudy (denisaboichuk):

-4 and 3